# On the Suitability of Reinforcement Fine-Tuning to Visual Tasks

### v1

@@ 3 Method
Normalized Length Reward. In VisualThinker [24], researchers observed that using a fixed-length reward of $+0.001$ per additional generated token led to an increase in response length without improving accuracy. The model exploited the reward by generating repetitive content rather than enhancing its reasoning process. To address this issue, w e apply sigmoid normalization to the length reward, stabilizing its value between 0 and 1, as follows:
@@
where $\lambda$ is a scaling factor set to 1 in our experiments, and $L_{0}$ is a predefined length fixed at $100$ across all experiments. Therefore, the length reward is kept within a reasonable range, as the diminishing marginal benefits help stabilize response length effectively.

### v2

@@ 3 Method
Normalized Length Reward. In VisualThinker [24], researchers observed that using a fixed-length reward of +0.001+0.001 per additional generated token led to an increase in response length without improving accuracy. The model exploited the reward by generating repetitive content rather than enhancing its reasoning process. To address this issue, w e apply sigmoid normalization to the length reward, stabilizing its value between 0 and 1, as follows:
@@
where $\lambda$ is a scaling factor set to 1 in our experiments, and $L_{0}$ is a predefined length fixed at 100100 across all experiments. Therefore, the length reward is kept within a reasonable range, as the diminishing marginal benefits help stabilize response length effectively.